PartI n an inertial frame of reference, a series of experiments is conducted. In each experiment, two or three forces are applied to an object. The magnitudes of these forces are given. No other forces are acting on the object. In which cases may the object possibly move at a constant velocity of 256 m/s? The forces applied are as follows: Check all that apply Hints Hint 1. Using the net force In an inertial frame of reference, moving at a constant velocity is only possible when the net force acting on the object iz zero. 2N: 2 N 200 N: 200 N 200 N: 201 N 2N: 2 N: 4 N 2 N: 2N: 2 N 2N: 2 N: 3 N 2N: 2 N: 5 N 200 N: 200 N: 5N

This is wrong.
Answer #1

N Net foyce 2-2 = ON obje ct opotite cie cne (achng in Net fore = 2o -2oo = 0 A 2N Nhe Net fra -2-2 = 4- 4oN So, ams rers are